OpenStudy (anonymous):

Use the substitution method to evalute the integral ∫((2x^4) +1))^4 x^3 dx. I managed to find u = 2x^4 + 1, then I found du = 8x^3 dx, and then du/8 = x^3 dx. I am totally baffled as to what the next step is?

OpenStudy (turingtest):

You are so close!\[\int(2x^4+1)^4x^3dx\]substitute what you have and the magic occurs\[u=2x^4+1\]\[\frac{du}8=x^3dx\]

OpenStudy (turingtest):

\[\int u^4(\frac{du}8)=\frac18\int u^4du\]

OpenStudy (turingtest):

@tbreez712 does this make sense to you?

OpenStudy (anonymous):

Sorry I was walking my dog. I'm a little confused as to what you did?

OpenStudy (turingtest):

\[\int(2x^4+1)^4x^3dx\]\[u=2x^4+1\]\[\frac{du}8=x^3dx\]

OpenStudy (anonymous):

yep yep

OpenStudy (turingtest):

so all we have is a substitution just make sure you recognize which part goes where putting a few more parentheses may help:\[\int(2x^4+1)^4(x^3dx)\]\[u=2x^4+1\]\[\frac{du}8=x^3dx\]so substituting in terms of u we get\[\int(u^4)(\frac{du}8)\]looking reasonable?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

we can take the constants out of the integral, so this is\[\frac18\int u^4du\]you should be able to integrate that what do you get?

OpenStudy (anonymous):

so aall you did there was rewrite du/8 aas 1/8? And if i try to integrate that I would get u5/5?

OpenStudy (turingtest):

I rewrote du/8 as du*(1/8) and took the 1/8 out of the integral sign and yes, the integral itself is (u^5)/5 don't forget to keep the constant though!

OpenStudy (anonymous):

sorry for sounding like a hillbilly, but is du the constant your talking about?

OpenStudy (turingtest):

no the constant is the number 1/8 we can take constants out of the integral sign it just makes it easier on the eyes\[\int2f(x)dx=2\int f(x)dx\]

OpenStudy (anonymous):

ohhh ! i get it!

OpenStudy (anonymous):

so 1/8 * u5/5?

OpenStudy (anonymous):

u5/40 substituted back ian so (2x4+1)^5 / 40?

OpenStudy (turingtest):

yep :) nice job!

OpenStudy (anonymous):

ahhh your awesome

OpenStudy (turingtest):

+C

OpenStudy (anonymous):

thanks for sticking through it with me!

OpenStudy (anonymous):

aoh yeah forgot the +c

OpenStudy (turingtest):

don't forget +C and you're very welcome